OpenStudy (anonymous):

You have $2,000 to invest, and want it to grow to $3,000 in two years. What interest rate would you need to find to make this possible?

OpenStudy (wolf1728):

The compound Interest formula is Total = Beginning Amount * (1 + rate)^years We have to solve this for rate log(1 + rate) = {log(total) -log(Principal)} ÷ Years log(1 + rate) = {log(3,000)-log(2,000)} ÷ 2 log(1 + rate) = { 3.4771212547 - 3.3010299957 } ÷ 2 log(1 + rate) = 0.176091259 ÷ 2 log(1 + rate) = 0.0880456295 1 + rate = 10^0.0880456295 1 + rate = 1.2247448713 rate = .2247448713 OR rate = 22.47448713% That is a VERY high interest rate because you want to increase your money by 150% in just two years.